1.Calcuate the pH of a solution with [H3O+] = 6.54 x 10-3 M

2.Calcuate the pOH of a solution with [H3O+] = 8.5 x 10-9 M
Answer #1

Note : [H3O+] represents the concentration of H+

Question 1

[H3O+] = 6.54×10-3 M

or

[H+] = 6.54×10-3 M

Using the formula

P​​​​​​H​​​​​ = -log[H​​​​​​+] = -log(6.5×10-3] = 2.187

P​​​​​​H = 2.187

Question 2

First of all we calculate P​​​​​​H

[H3O+] = [H+] = 8.5×10-9

P​​​​​​H = -log [H+] = -log(8.5×10-9) = 8.070

P​​​​​​H = 8.070

now using the formula

P​​​​​​H​​​​​ + P​​​​​​OH = 14

8.070 + P​​​​​​OH = 14

P​​​​​​OH = 5.93